OpenStudy (anonymous):

Can someone help me solve this scientific notation calculation problem? (3.4 x 10^-1)^3

OpenStudy (anonymous):

3.4*10^-1=3.4*0.1=0.34 o.34^3=0.0393

OpenStudy (anonymous):

Thanks!